# Dynamical Dark Energy Beyond Planck? Constraints from multiple CMB probes, DESI BAO and Type-Ia Supernovae

### v1

@@ III.3 Joint Constraints from Planck, ACT and SPT
Note that I retain the Planck temperature and polarization measurements at large angular scales while cutting the high-multipole measurements at $\ell=650$ to avoid double-counting the same portion of the sky already measured by ACT and SPT data. [^3] This is the same approach followed by the ACT collaboration, see e.g., Ref Aiola et al. (2020). This reduced information in the Planck dataset allows me to combine this likelihood with the full ACT-DR4 and ACT-DR6 temperature, polarization, and lensing likelihoods, as well as with the full SPT temperature and polarization likelihoods.
The numerical constraints on the cosmological parameters obtained by combining Planck-650, ACT, SPT, DESI BAO measurements, and type Ia SN from either the PP or DESy5 catalogues are presented in Appendix A (last columns of Tabs. 5, 6, 7, 8) and summarized in Fig. 3 (see also Fig. 6 and Fig. 7 in Appendix A) . As seen in the top panel of the figure, when considering the combinations ACT+Planck-650+DESI+PP or ACT+Planck-650+DESI+DESy5, the preference for DDE is well confirmed in both cases. In the former case, the constraints on the DE parameters are $w_{0}=-0.82^{+0.13}_{-0.12}$ and $w_{a}=-0.74^{+0.50}_{-0.57}$, while for the latter dataset, I get $w_{0}=-0.73\pm 0.13$ and $w_{a}=-1.01^{+0.55}_{-0.63}$ (all at 95% CL). Therefore, the preference is strengthened compared to the same data combinations where Planck-650 was replaced by WMAP, while appearing very similar to Planck+DESI+PP or Planck+DESI+DESy5, respectively. This confirms the overall conclusion that part of the preference for DDE is driven by large-scale temperature and polarization measurements at $\ell\lesssim 30$ (specifically E-mode polarization measurements), as this portion of information is actually missing in WMAP. On the other hand, from the bottom panel of Fig. 3, we see that no clear preference for DDE is observed in the SPT-based results. In the case of SPT+Planck-650+DESI+PP, the cosmological constant falls well within the 95% CL and the constraints are $w_{0}=-0.86\pm 0.19$ and $w_{a}=-0.57\pm 0.89$ at 95%CL. For SPT+Planck-650+DESI+DESy5, the cosmological constant falls at the boundary of the contours as I get $w_{0}=-0.74\pm 0.21$ and $w_{a}=-1.0\pm 1.0$, always at 95%CL. Overall, the interpretation of these results remains challenging due to the much larger uncertainties.
@@ IV Conclusion
The DESI collaboration has recently released BAO measurements that, when combined with Planck CMB data and Type Ia Supernovae from either the Pantheon-plus or DESy5 surveys, produce a preference for Dynamical Dark Energy at a significance ranging between 3.5 and 3.9 standard deviations. In this work, I investigate the role of Planck CMB measurements in favoring this preference and test whether and to what extent the shift toward evolving Dark Energy is supported by other CMB experiments.
@@
Specifically, in Sec. III.1, I first focus Planck-independent combinations of CMB experiments analyzing the Atacama Cosmology Telescope and South Pole Telescope measurements of temperature, polarization, and lensing spectra at small scales, eventually combined with WMAP 9-year observations to include information at large angular scales. My analysis reveals that CMB experiments other than Planck tend to reduce the preference for evolving Dark Energy. This conclusion is evident from the numerical results on the dark energy parameters $w_{0}$ and $w_{a}$ summarized in Tab. 2, as well as from Fig. 1, which shows the marginalized probability contours at 68% and 95% confidence levels in the $w_{0}$-$w_{a}$ plane. As seen in Fig. 1 and Tab. 2, when the Atacama Cosmology Telescope temperature, polarization, and lensing data (either alone or combined with WMAP 9-year observations) are used alongside DESI BAO and supernova measurements, they achieve a constraining power comparable to that of Planck. However, the preference for Dynamical Dark Energy is significantly reduced when focusing on the combination involving ACT(+WMAP), DESI BAO, and Pantheon-plus supernovae. In this case, the cosmological constant falls within the 95% confidence level contours. A preference for Dynamical Dark Energy reemerges when Pantheon-plus is replaced with DESy5 supernova data, though it remains less pronounced compared to results based on Planck CMB measurements. This reduction in the preference for evolving Dark Energy is confirmed when comparing the improvement in the fit over $\Lambda$CDM. This information is detailed in the last column of Tab. 2, where I report the best-fit $\Delta\chi^{2}=\chi^{2}_{\text{CPL}}-\chi^{2}_{\Lambda\text{CDM}}$ for each combination of datasets analyzed in my work. The improvement in the fit is always reduced in ACT-based results compared to those based on Planck data. The interpretation of the results becomes somewhat more challenging when extending the analysis to the SPT data. In this case, the larger uncertainties obtained in inferring cosmological parameters present both a challenge and a concern, introducing significant correlations among the dark energy parameters $w_{0}$ and $w_{a}$ and other crucial parameters characterizing the late time dynamics such as $H_{0}$ and $\Omega_{m}$, see also Fig. 4 and Fig. 5 in Appendix A. Having that said, as seen from Fig. 1, the preference for Dynamical Dark Energy is essentially lost, regardless of the SN datasets used in the analysis. This is also reflected in the modest improvement in $\Delta\chi^{2}$.
@@
Given these results, as a second step, I further investigate the role of Planck data. In Sec. III.2, I argue that the subset of Planck measurements strengthening this preference are the temperature and polarization measurements at large angular scales $\ell\lesssim 30$, especially the E-mode polarization measurements. This is somewhat expected, as the most significant impact on the CMB angular power spectrum due to Dark Energy dynamics arises from changes in the amplitude of the Integrated Sachs-Wolfe plateau affecting these scales. As seen in Fig. 2, cutting down information at $\ell\lesssim 30$ from the Planck likelihoods, the cosmological constant value falls within the 95% CL contours for the constraints based on Pantheon-plus supernovae due to a significant shift in the central value of $w_{a}$. On the other hand, a preference for evolving Dark Energy persists in the results based on DESy5 supernovae, although to a lesser extent. To further validate my conclusions regarding the importance of Planck’s temperature and polarization measurements at large angular scales, in Sec. III.3, I perform a joint analysis of Planck large-scale temperature and polarization data (which I argued strengthen the preference) with the Atacama Cosmology Telescope and South Pole Telescope small-scale measurements (which, when analyzed on their own or in combination with WMAP, reduce the preference). The joint analysis supports the interpretation that part of the shift toward Dynamical Dark Energy is reinforced by Planck’s large-scale temperature and polarization anisotropy measurements. Indeed, as seen in Fig. 3, when Planck’s large-scale data are analyzed in combination with the Atacama Cosmology Telescope likelihoods, the preference for evolving Dark Energy is significantly strengthened compared to the case without Planck. Interestingly, however, no preference is observed when combining Planck’s large-scale data with the South Pole Telescope likelihoods, although the larger uncertainties in the parameter space remain a significant limitation.
@@
  • No convincing preference is seen in any combinations involving South Pole Telescope data, whether considering Pantheon-plus, DESy5, or including Planck/WMAP large-scale temperature and polarization measurements.
When it comes to interpreting these findings, they present a double-edged sword. On the one hand, one could argue that retaining all information from large to small scales across different CMB experiments (see, e.g., Planck-650+ACT-based results) still hints at a preference for Dynamical Dark Energy, which could be interpreted as a resilience of the signal. On the other hand, it remains true that this preference is primarily reinforced by Planck’s large-scale data and the DESy5 supernovae. The former have consistently driven mild anomalies and deviations from $\Lambda$CDM predictions Escamilla et al. (2024b); Giarè et al. (2024c); Ben-Dayan et al. (2024), while the latter have recently been subject to debate due to potential calibration issues Efstathiou (2024). Additionally, the presence of unexplained outliers in the preliminary DESI BAO measurements could also contribute to the overall shift toward Dynamical Dark Energy Wang et al. (2024d); Colgáin et al. (2024); Naredo-Tuero et al. (2024). Wrapping everything up, for those of us adopting a more conservative stance, these findings may warrant additional caution in assessing the robustness of the DESI BAO and supernova preference for Dynamical Dark Energy. Further tests and future data – including upcoming CMB observations and future BAO data releases from DESI – will be invaluable in providing a definitive answer.
@@ Acknowledgements.
I thank Eleonora Di Valentino, Alessandro Melchiorri, and Elsa M. Teixeira for interesting discussions and suggestions on a preliminary version of this work. I acknowledge support from the Lancaster–Sheffield Consortium for Fundamental Physics through the Science and Technology Facilities Council (STFC) grant ST/X000621/1. This article is based upon work from COST Action CA21136 Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse) supported by COST (European Cooperation in Science and Technology). I acknowledge IT Services at The University of Sheffield for the provision of services for High Performance Computing.
@@ Appendix A Numerical constraints on cosmological parameters and contour plots
$w_{0}$ $-0.826\pm 0.062\,(-0.83^{+0.12}_{-0.12})$ $-0.827\pm 0.063\,(-0.83^{+0.13}_{-0.12})$ $-0.847\pm 0.064\,(-0.85^{+0.13}_{-0.12})$ $-0.850\pm 0.065\,(-0.85^{+0.13}_{-0.13})$
$w_{a}$ $-0.72^{+0.29}_{-0.25}\,(-0.72^{+0.50}_{-0.56})$ $-0.72^{+0.30}_{-0.24}\,(-0.72^{+0.51}_{-0.58})$ $-0.56^{+0.32}_{-0.26}\,(-0.56^{+0.55}_{-0.58})$ $-0.55^{+0.32}_{-0.27}\,(-0.55^{+0.56}_{-0.61})$
$\Omega_{\mathrm{b}}h^{2}$ $0.02243\pm 0.00014\,(0.02243^{+0.00027}_{-0.00027})$ $0.02241\pm 0.00014\,(0.02241^{+0.00027}_{-0.00027})$ $0.02250\pm 0.00015\,(0.02250^{+0.00031}_{-0.00030})$ $0.02250\pm 0.00016\,(0.02250^{+0.00031}_{-0.00030})$
$\Omega_{\mathrm{c}}h^{2}$ $0.1194\pm 0.0010\,(0.1194^{+0.0020}_{-0.0020})$ $0.1196\pm 0.0010\,(0.1196^{+0.0020}_{-0.0020})$ $0.1185\pm 0.0013\,(0.1185^{+0.0025}_{-0.0026})$ $0.1185\pm 0.0014\,(0.1185^{+0.0027}_{-0.0027})$
$\tau_{\mathrm{reio}}$ $0.0538\pm 0.0073\,(0.054^{+0.015}_{-0.014})$ $0.0537\pm 0.0073\,(0.054^{+0.015}_{-0.014})$ $0.072\pm 0.018\,(0.072^{+0.035}_{-0.034})$ $0.075\pm 0.019\,(0.075^{+0.038}_{-0.036})$
$H_{0}$ $68.00\pm 0.71\,(68.0^{+1.4}_{-1.4})$ $67.99\pm 0.72\,(68.0^{+1.4}_{-1.4})$ $67.91\pm 0.73\,(67.9^{+1.4}_{-1.4})$ $67.85\pm 0.73\,(67.9^{+1.5}_{-1.4})$
$n_{\mathrm{s}}$ $0.9669\pm 0.0038\,(0.9669^{+0.0075}_{-0.0075})$ $0.9654\pm 0.0039\,(0.9654^{+0.0076}_{-0.0077})$ $0.9697\pm 0.0047\,(0.9697^{+0.0092}_{-0.0091})$ $0.9688\pm 0.0048\,(0.9688^{+0.0098}_{-0.0092})$
$\log(10^{10}A_{\mathrm{s}})$ $3.042\pm 0.014\,(3.042^{+0.029}_{-0.027})$ $3.042\pm 0.014\,(3.042^{+0.029}_{-0.029})$ $3.076\pm 0.033\,(3.076^{+0.065}_{-0.063})$ $3.081\pm 0.035\,(3.081^{+0.069}_{-0.066})$
@@
$w_{0}$ $-0.736\pm 0.068\,(-0.74^{+0.14}_{-0.13})$ $-0.735\pm 0.069\,(-0.73^{+0.14}_{-0.13})$ $-0.757\pm 0.070\,(-0.76^{+0.14}_{-0.13})$ $-0.758^{+0.067}_{-0.075}\,(-0.76^{+0.14}_{-0.13})$
$w_{a}$ $-0.99^{+0.33}_{-0.28}\,(-0.99^{+0.57}_{-0.64})$ $-1.00^{+0.33}_{-0.28}\,(-1.00^{+0.57}_{-0.65})$ $-0.84^{+0.36}_{-0.30}\,(-0.84^{+0.63}_{-0.66})$ $-0.84^{+0.36}_{-0.30}\,(-0.84^{+0.63}_{-0.67})$
$\Omega_{\mathrm{b}}h^{2}$ $0.02241\pm 0.00014\,(0.02241^{+0.00027}_{-0.00027})$ $0.02240\pm 0.00014\,(0.02240^{+0.00027}_{-0.00027})$ $0.02247\pm 0.00016\,(0.02247^{+0.00031}_{-0.00031})$ $0.02247\pm 0.00015\,(0.02247^{+0.00031}_{-0.00030})$
$\Omega_{\mathrm{c}}h^{2}$ $0.1196\pm 0.0010\,(0.1196^{+0.0020}_{-0.0020})$ $0.1198\pm 0.0010\,(0.1198^{+0.0019}_{-0.0020})$ $0.1188\pm 0.0013\,(0.1188^{+0.0026}_{-0.0026})$ $0.1189\pm 0.0013\,(0.1189^{+0.0026}_{-0.0026})$
$\tau_{\mathrm{reio}}$ $0.0527\pm 0.0074\,(0.053^{+0.015}_{-0.015})$ $0.0529\pm 0.0073\,(0.053^{+0.014}_{-0.015})$ $0.069\pm 0.018\,(0.069^{+0.035}_{-0.035})$ $0.070\pm 0.018\,(0.070^{+0.036}_{-0.035})$
$H_{0}$ $67.25\pm 0.66\,(67.2^{+1.3}_{-1.3})$ $67.23\pm 0.65\,(67.2^{+1.3}_{-1.3})$ $67.20\pm 0.66\,(67.2^{+1.3}_{-1.3})$ $67.18\pm 0.67\,(67.2^{+1.3}_{-1.3})$
$n_{\mathrm{s}}$ $0.9663\pm 0.0038\,(0.9663^{+0.0075}_{-0.0075})$ $0.9648\pm 0.0039\,(0.9648^{+0.0078}_{-0.0076})$ $0.9689\pm 0.0046\,(0.9689^{+0.0092}_{-0.0091})$ $0.9675\pm 0.0047\,(0.9675^{+0.0094}_{-0.0092})$
$\log(10^{10}A_{\mathrm{s}})$ $3.040\pm 0.015\,(3.040^{+0.029}_{-0.029})$ $3.041\pm 0.014\,(3.041^{+0.028}_{-0.029})$ $3.069\pm 0.033\,(3.069^{+0.065}_{-0.064})$ $3.072\pm 0.033\,(3.072^{+0.066}_{-0.064})$
@@
$w_{0}$ $-0.854\pm 0.065\,(-0.85^{+0.13}_{-0.12})$ $-0.850\pm 0.064\,(-0.85^{+0.13}_{-0.12})$ $-0.823\pm 0.064\,(-0.82^{+0.13}_{-0.12})$
$w_{a}$ $-0.55^{+0.33}_{-0.28}\,(-0.55^{+0.55}_{-0.62})$ $-0.55^{+0.31}_{-0.26}\,(-0.55^{+0.56}_{-0.58})$ $-0.74^{+0.29}_{-0.25}\,(-0.74^{+0.50}_{-0.57})$
$\Omega_{\mathrm{b}}h^{2}$ $0.02161\pm 0.00029\,(0.02161^{+0.00057}_{-0.00058})$ $0.02241\pm 0.00018\,(0.02241^{+0.00035}_{-0.00035})$ $0.02237\pm 0.00012\,(0.02237^{+0.00023}_{-0.00022})$
$\Omega_{\mathrm{c}}h^{2}$ $0.1189\pm 0.0016\,(0.1189^{+0.0031}_{-0.0030})$ $0.1191\pm 0.0014\,(0.1191^{+0.0027}_{-0.0028})$ $0.11974\pm 0.00091\,(0.1197^{+0.0018}_{-0.0018})$
$\tau_{\mathrm{reio}}$ $0.066\pm 0.013\,(0.066^{+0.025}_{-0.026})$ $0.063\pm 0.013\,(0.063^{+0.025}_{-0.024})$ $0.0544\pm 0.0070\,(0.054^{+0.014}_{-0.013})$
$H_{0}$ $67.50\pm 0.74\,(67.5^{+1.5}_{-1.4})$ $67.86\pm 0.73\,(67.9^{+1.4}_{-1.4})$ $67.98\pm 0.73\,(68.0^{+1.4}_{-1.4})$
$n_{\mathrm{s}}$ $0.998\pm 0.012\,(0.998^{+0.023}_{-0.023})$ $0.9748\pm 0.0044\,(0.9748^{+0.0088}_{-0.0086})$ $0.9692\pm 0.0034\,(0.9692^{+0.0068}_{-0.0066})$
$\log(10^{10}A_{\mathrm{s}})$ $3.063\pm 0.024\,(3.063^{+0.046}_{-0.047})$ $3.067\pm 0.022\,(3.067^{+0.045}_{-0.043})$ $3.050\pm 0.013\,(3.050^{+0.026}_{-0.025})$

### v2

@@ III.3 Joint Constraints from Planck, ACT and SPT
Note that I retain the Planck temperature and polarization measurements at large angular scales while cutting the high-multipole measurements at $\ell=650$ to avoid double-counting the same portion of the sky already measured by ACT and SPT data. [^5] This is the same approach followed by the ACT collaboration, see e.g., Ref Aiola et al. (2020); Louis et al. (2025). This reduced information in the Planck dataset allows me to combine this likelihood with the full temperature, polarization, and lensing ACT-DR6 likelihoods, as well as with the full SPT temperature and polarization likelihoods.
The numerical constraints on the cosmological parameters obtained by combining Planck-650, ACT, SPT, DESI BAO measurements, and type Ia SN from either the PP or DESy5 catalogues are presented in Appendix A (last columns of Tabs. 5, 6, 7, 8) and summarized in Fig. 3 (see also Fig. 6 and Fig. 7 in Appendix A) . As seen in the top panel of the figure, when considering the combinations ACT+Planck-650+DESI+PP or ACT+Planck-650+DESI+DESy5, the preference for DDE is well confirmed in both cases. In the former case, the constraints on the DE parameters are $w_{0}=-0.845\pm 0.054$ and $w_{a}=-0.55^{+0.21}_{-0.19}$, while for the latter dataset, I get $w_{0}=-0.759\pm 0.057$ and $w_{a}=-0.79^{+0.23}_{-0.20}$ (all at 68% CL). Therefore, the preference is strengthened compared to the same data combinations where Planck-650 was replaced by WMAP, while appearing very similar to Planck+DESI+PP or Planck+DESI+DESy5, respectively. This confirms the overall conclusion that part of the preference for DDE is driven by large-scale temperature and polarization measurements at $\ell\lesssim 30$ (specifically E-mode polarization measurements, see also Appendix B), as this portion of information is actually missing in WMAP. On the other hand, from the bottom panel of Fig. 3, no clear preference for DDE is observed in the SPT+PP-based results. In the case of SPT+Planck-650+DESI+PP, the cosmological constant lies within the 95% CL, with constraints $w_{0}=-0.860\pm 0.056$ and $w_{a}=-0.46^{+0.23}_{-0.21}$ at 68% CL. For SPT+Planck-650+DESI+DESy5, a preference for dynamical dark energy is again observed (although weakened), with $w_{0}=-0.774\pm 0.059$ and $w_{a}=-0.71^{+0.26}_{-0.22}$ at 68% CL.
@@ IV Conclusion
The DESI collaboration has recently released BAO measurements based on its second year of observations that, when combined with Planck CMB data and Type Ia Supernovae from either the Pantheon-plus or DESy5 surveys, produce a preference for Dynamical Dark Energy at a significance ranging between 2.8 and 4.2 standard deviations. In this work, I investigate the role of Planck CMB measurements in favoring this preference and test whether and to what extent the shift toward evolving Dark Energy is supported by other CMB experiments.
@@
Specifically, in Sec. III.1, I first focus Planck-independent combinations of CMB experiments analyzing the latest Atacama Cosmology Telescope and South Pole Telescope measurements of temperature, polarization, and lensing spectra at small scales, eventually combined with WMAP 9-year observations to include information at large angular scales. My analysis reveals that CMB experiments other than Planck tend to reduce the preference for evolving Dark Energy. This conclusion is evident from the numerical results on the dark energy parameters $w_{0}$ and $w_{a}$ summarized in Tab. 2, as well as from Fig. 1, which shows the marginalized probability contours at 68% and 95% confidence levels in the $w_{0}$-$w_{a}$ plane. As seen in Fig. 1 and Tab. 2, when the Atacama Cosmology Telescope temperature, polarization, and lensing data (either alone or combined with WMAP 9-year observations) are used alongside DESI BAO and supernova measurements, they achieve a constraining power comparable to that of Planck. However, the preference for Dynamical Dark Energy is significantly reduced when focusing on the combination involving ACT(+WMAP), DESI BAO, and Pantheon-plus supernovae. In this case, the cosmological constant falls close to the 95% confidence level contours. A preference for Dynamical Dark Energy reemerges when Pantheon-plus is replaced with DESy5 supernova data, though it remains less pronounced compared to results based on Planck CMB measurements. This reduction in the preference for evolving Dark Energy is confirmed when comparing the improvement in the fit over $\Lambda$CDM. This information is detailed in the last column of Tab. 2, where I report the best-fit $\Delta\chi^{2}=\chi^{2}_{\text{CPL}}-\chi^{2}_{\Lambda\text{CDM}}$ for each combination of datasets analyzed in my work. The improvement in the fit is always reduced in ACT-based results compared to those based on Planck data. The interpretation of the results becomes somewhat more challenging when extending the analysis to the SPT data. In this case, the larger uncertainties obtained in inferring cosmological parameters present both a challenge and a concern, introducing significant correlations among the dark energy parameters $w_{0}$ and $w_{a}$ and other crucial parameters characterizing the late time dynamics such as $H_{0}$ and $\Omega_{m}$, see also Fig. 4 and Fig. 5 in Appendix A. That said, as seen in Fig. 1, the preference for Dynamical Dark Energy is essentially lost when SPT is combined with Pantheon-plus SN and is significancy reduced for combinations involving DESy5. This is also reflected in the behavior of $\Delta\chi^{2}$.
@@
Given these results, as a second step, I further investigate the role of Planck data. In Sec. III.2, I argue that the subset of Planck measurements strengthening this preference are the temperature and polarization measurements at large angular scales $\ell\lesssim 30$, especially the E-mode polarization measurements. This is somewhat expected, as the most significant impact of Dark Energy dynamics on the CMB angular power spectrum typically arises from changes in the amplitude of the Integrated Sachs-Wolfe plateau, which affect these scales in the TT spectrum, or from correlations between Dark Energy parameters and other cosmological parameters such as $\tau$, $A_{s}$, and $\Omega_{m}$, which can also be sensitive to arge scale measurements (see, e.g., Ref. Giarè et al. (2024c) and Appendix B). As seen in Fig. 2, cutting down information at $\ell\lesssim 30$ from the Planck likelihoods, the cosmological constant value falls within the 95% CL contours for the constraints based on Pantheon-plus supernovae due to a significant shift in the central value of $w_{a}$. On the other hand, a preference for evolving Dark Energy persists in the results based on DESy5 supernovae, although to a lesser extent. To further validate my conclusions regarding the importance of Planck’s temperature and polarization measurements at large angular scales, in Sec. III.3, I perform a joint analysis of Planck large-scale temperature and polarization data (which I argued strengthen the preference) with the Atacama Cosmology Telescope and South Pole Telescope small-scale measurements (which, when analyzed on their own or in combination with WMAP, reduce the preference). The joint analysis supports the interpretation that part of the shift toward Dynamical Dark Energy is reinforced by Planck’s large-scale temperature and polarization anisotropy measurements. Indeed, as seen in Fig. 3, when Planck’s large-scale data are analyzed in combination with the Atacama Cosmology Telescope likelihoods, the preference for evolving Dark Energy is strengthened compared to the case without Planck. Interestingly, no strong preference is observed when combining Planck’s large-scale data with the South Pole Telescope likelihoods alongside Pantheon-plus SN.
@@
  • No convincing preference is seen in any combinations involving South Pole Telescope data and Pantheon-plus, while it is reduced in combinations involving South Pole Telescope data and DESy5 SN (both including or excluding Planck/WMAP large-scale temperature and polarization measurements).
When it comes to interpreting these findings, they present a double-edged sword. On the one hand, one could argue that retaining all information from large to small scales across different CMB experiments (see, e.g., Planck-650+ACT-based results) still hints at a preference for Dynamical Dark Energy, which could be interpreted as a resilience of the signal Giarè et al. (2025). On the other hand, it remains true that this preference is primarily reinforced by Planck’s large-scale data and the DESy5 supernovae. The former have consistently driven mild anomalies and deviations from $\Lambda$CDM predictions Escamilla et al. (2024b); Giarè et al. (2024c); Ben-Dayan et al. (2025), while the latter have recently been subject to debate due to potential calibration issues Efstathiou (2025). Additionally, the presence of unexplained outliers in the preliminary DESI BAO measurements could also contribute to the overall shift toward Dynamical Dark Energy Wang et al. (2024c); Colgáin et al. (2024a); Naredo-Tuero et al. (2024). Wrapping everything up, for those of us adopting a more conservative stance, these findings may warrant additional caution in assessing the robustness of the DESI BAO and supernova preference for Dynamical Dark Energy. Further tests and future data – including upcoming CMB observations and future BAO data releases from DESI – will be invaluable in providing a definitive answer.
@@ Acknowledgements.
I thank Eleonora Di Valentino, Alessandro Melchiorri, and Elsa M. Teixeira for interesting discussions and suggestions on a preliminary version of this work. I also thank the referee for the valuable suggestions that led to the additional tests presented in Appendix B, which offer further insight into the interpretation of the results. I acknowledge support from the Lancaster–Sheffield Consortium for Fundamental Physics through the Science and Technology Facilities Council (STFC) grant ST/X000621/1. This article is based upon work from COST Action CA21136 Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse) supported by COST (European Cooperation in Science and Technology). I acknowledge IT Services at The University of Sheffield for the provision of services for High Performance Computing.
@@ Appendix A Numerical constraints on cosmological parameters and contour plots
$w_{0}$ $-0.846\pm 0.054\,(-0.85^{+0.11}_{-0.11})$ $-0.844\pm 0.054\,(-0.84^{+0.11}_{-0.11})$ $-0.867\pm 0.055\,(-0.87^{+0.11}_{-0.11})$ $-0.868\pm 0.056\,(-0.87^{+0.11}_{-0.11})$
$w_{a}$ $-0.56^{+0.22}_{-0.19}\,(-0.56^{+0.39}_{-0.42})$ $-0.58^{+0.22}_{-0.19}\,(-0.58^{+0.39}_{-0.42})$ $-0.41^{+0.24}_{-0.20}\,(-0.41^{+0.42}_{-0.43})$ $-0.41^{+0.23}_{-0.21}\,(-0.41^{+0.42}_{-0.45})$
$\Omega_{\mathrm{b}}h^{2}$ $0.02245\pm 0.00013\,(0.02245^{+0.00026}_{-0.00026})$ $0.02245\pm 0.00014\,(0.02245^{+0.00027}_{-0.00027})$ $0.02253\pm 0.00014\,(0.02253^{+0.00029}_{-0.00028})$ $0.02254\pm 0.00015\,(0.02254^{+0.00030}_{-0.00029})$
$\Omega_{\mathrm{c}}h^{2}$ $0.11901\pm 0.00087\,(0.1190^{+0.0017}_{-0.0017})$ $0.11917\pm 0.00090\,(0.1192^{+0.0017}_{-0.0018})$ $0.1180\pm 0.0011\,(0.1180^{+0.0022}_{-0.0022})$ $0.1181\pm 0.0011\,(0.1181^{+0.0022}_{-0.0023})$
$\tau_{\mathrm{reio}}$ $0.0555\pm 0.0072\,(0.055^{+0.015}_{-0.014})$ $0.0553\pm 0.0074\,(0.055^{+0.015}_{-0.014})$ $0.078\pm 0.017\,(0.078^{+0.033}_{-0.033})$ $0.080\pm 0.017\,(0.080^{+0.033}_{-0.033})$
$H_{0}$ $67.65\pm 0.60\,(67.6^{+1.2}_{-1.2})$ $67.65\pm 0.59\,(67.7^{+1.2}_{-1.1})$ $67.60\pm 0.59\,(67.6^{+1.2}_{-1.2})$ $67.60\pm 0.59\,(67.6^{+1.2}_{-1.1})$
$n_{\mathrm{s}}$ $0.9679\pm 0.0036\,(0.9679^{+0.0071}_{-0.0071})$ $0.9664\pm 0.0037\,(0.9664^{+0.0072}_{-0.0072})$ $0.9710\pm 0.0043\,(0.9710^{+0.0086}_{-0.0082})$ $0.9699\pm 0.0043\,(0.9699^{+0.0087}_{-0.0085})$
$\log(10^{10}A_{\mathrm{s}})$ $3.045\pm 0.014\,(3.045^{+0.028}_{-0.028})$ $3.045\pm 0.015\,(3.045^{+0.029}_{-0.028})$ $3.087\pm 0.031\,(3.087^{+0.060}_{-0.061})$ $3.090\pm 0.032\,(3.090^{+0.061}_{-0.062})$
@@
$w_{0}$ $-0.760\pm 0.057\,(-0.76^{+0.11}_{-0.11})$ $-0.759\pm 0.056\,(-0.76^{+0.11}_{-0.11})$ $-0.783\pm 0.058\,(-0.78^{+0.12}_{-0.11})$ $-0.782\pm 0.058\,(-0.78^{+0.12}_{-0.11})$
$w_{a}$ $-0.80\pm 0.22\,(-0.80^{+0.42}_{-0.45})$ $-0.81^{+0.23}_{-0.20}\,(-0.81^{+0.42}_{-0.45})$ $-0.66^{+0.26}_{-0.22}\,(-0.66^{+0.46}_{-0.48})$ $-0.67^{+0.25}_{-0.22}\,(-0.67^{+0.46}_{-0.50})$
$\Omega_{\mathrm{b}}h^{2}$ $0.02244\pm 0.00013\,(0.02244^{+0.00027}_{-0.00026})$ $0.02244\pm 0.00013\,(0.02244^{+0.00026}_{-0.00026})$ $0.02251\pm 0.00015\,(0.02251^{+0.00028}_{-0.00028})$ $0.02251\pm 0.00015\,(0.02251^{+0.00029}_{-0.00028})$
$\Omega_{\mathrm{c}}h^{2}$ $0.11925\pm 0.00090\,(0.1192^{+0.0017}_{-0.0018})$ $0.11936\pm 0.00090\,(0.1194^{+0.0017}_{-0.0018})$ $0.1184\pm 0.0011\,(0.1184^{+0.0022}_{-0.0023})$ $0.1184\pm 0.0011\,(0.1184^{+0.0022}_{-0.0022})$
$\tau_{\mathrm{reio}}$ $0.0544\pm 0.0072\,(0.054^{+0.015}_{-0.014})$ $0.0546\pm 0.0072\,(0.055^{+0.014}_{-0.014})$ $0.074\pm 0.017\,(0.074^{+0.033}_{-0.033})$ $0.076\pm 0.017\,(0.076^{+0.033}_{-0.033})$
$H_{0}$ $66.88\pm 0.56\,(66.9^{+1.1}_{-1.1})$ $66.88\pm 0.56\,(66.9^{+1.1}_{-1.1})$ $66.87\pm 0.56\,(66.9^{+1.1}_{-1.1})$ $66.86\pm 0.56\,(66.9^{+1.1}_{-1.1})$
$n_{\mathrm{s}}$ $0.9673\pm 0.0036\,(0.9673^{+0.0070}_{-0.0071})$ $0.9660\pm 0.0037\,(0.9660^{+0.0073}_{-0.0071})$ $0.9699\pm 0.0043\,(0.9699^{+0.0085}_{-0.0085})$ $0.9689\pm 0.0044\,(0.9689^{+0.0086}_{-0.0085})$
$\log(10^{10}A_{\mathrm{s}})$ $3.043\pm 0.014\,(3.043^{+0.029}_{-0.028})$ $3.044\pm 0.014\,(3.044^{+0.028}_{-0.028})$ $3.079\pm 0.031\,(3.079^{+0.061}_{-0.061})$ $3.082\pm 0.031\,(3.082^{+0.061}_{-0.060})$
@@
$w_{0}$ $-0.852^{+0.052}_{-0.058}\,(-0.85^{+0.11}_{-0.11})$ $-0.860\pm 0.053\,(-0.86^{+0.11}_{-0.10})$ $-0.845\pm 0.054\,(-0.85^{+0.11}_{-0.11})$
$w_{a}$ $-0.52^{+0.24}_{-0.20}\,(-0.52^{+0.41}_{-0.47})$ $-0.46^{+0.22}_{-0.20}\,(-0.46^{+0.38}_{-0.43})$ $-0.55^{+0.21}_{-0.19}\,(-0.55^{+0.38}_{-0.41})$
$\Omega_{\mathrm{b}}h^{2}$ $0.02260\pm 0.00016\,(0.02260^{+0.00031}_{-0.00032})$ $0.02266\pm 0.00012\,(0.02266^{+0.00024}_{-0.00024})$ $0.02252\pm 0.00011\,(0.02252^{+0.00021}_{-0.00022})$
$\Omega_{\mathrm{c}}h^{2}$ $0.1188\pm 0.0012\,(0.1188^{+0.0023}_{-0.0024})$ $0.1184\pm 0.0011\,(0.1184^{+0.0022}_{-0.0022})$ $0.11874\pm 0.00085\,(0.1187^{+0.0016}_{-0.0017})$
$\tau_{\mathrm{reio}}$ $0.072\pm 0.012\,(0.072^{+0.023}_{-0.022})$ $0.070\pm 0.011\,(0.070^{+0.022}_{-0.021})$ $0.0569\pm 0.0071\,(0.057^{+0.014}_{-0.014})$
$H_{0}$ $67.67\pm 0.60\,(67.7^{+1.2}_{-1.2})$ $67.69\pm 0.61\,(67.7^{+1.2}_{-1.2})$ $67.64\pm 0.60\,(67.6^{+1.2}_{-1.2})$
$n_{\mathrm{s}}$ $0.9763\pm 0.0068\,(0.976^{+0.013}_{-0.013})$ $0.9721\pm 0.0042\,(0.9721^{+0.0083}_{-0.0082})$ $0.9726\pm 0.0032\,(0.9726^{+0.0062}_{-0.0066})$
$\log(10^{10}A_{\mathrm{s}})$ $3.072\pm 0.021\,(3.072^{+0.041}_{-0.041})$ $3.074\pm 0.020\,(3.074^{+0.040}_{-0.039})$ $3.048\pm 0.013\,(3.048^{+0.025}_{-0.025})$